OpenStudy (anonymous):

Need some help on this too! In your lab, a substance's temperature has been observed to follow the function T(x) = (x + 5)3 + 7. The turning point of the graph is where the substance changes from a liquid to a solid. Explain to your fellow scientists how to find the turning point of this function, using complete sentences.

OpenStudy (yttrium):

Find it's roots. That will tell you where the turning points are.

OpenStudy (anonymous):

How do I find its roots?

OpenStudy (yttrium):

Let T(x) = 0 Then substitute in the equation. 0 = (x+5)^3 +7 0 = x+5 + cuberoot(7) x = -5-cureboor(7) approximately 6.913 That is the turning point.

OpenStudy (yttrium):

Oops put negative sign on 6.913

OpenStudy (anonymous):

Thank you! Is that all?

OpenStudy (yttrium):

Yup.

OpenStudy (anonymous):

could you help with one more?

OpenStudy (yttrium):

Just tell them that to find the turning poin(s)t, you have to find the roots of the equation. Then elaborate the method in finding the roots. :)

OpenStudy (yttrium):

Le try

OpenStudy (yttrium):

But better close this one and post that as another question

OpenStudy (anonymous):

got it!